OpenStudy (anonymous):

find the circumference. angle is 135 and arc is 9.8

OpenStudy (anonymous):

The circumference is the entire length around the circle. So if 135 degrees corresponds to 9.8, then 135/360 = 9.8/C C = (9.8*360)/135 = 26.13333

OpenStudy (anonymous):

thank you